OpenStudy (anonymous):

What is the solution of the following DIFFERENTIAL EQUATION for the initial value problem: y(0) = -2 Eqn: y' + (tanx)y = 2(cosx)^2

OpenStudy (anonymous):

\[\frac{dy}{dx} + tanx(y) = 2\cos^2x\]

OpenStudy (anonymous):

@experimentX

OpenStudy (anonymous):

@apoorvk

OpenStudy (anonymous):

@Callisto

OpenStudy (apoorvk):

you mean tanx multiplied by y - right?

OpenStudy (anonymous):

Yes

OpenStudy (apoorvk):

Use the funda of Integrating factor - know it?

OpenStudy (anonymous):

I am trying to get the form: \[f(x)dx = f(y)dy\]

OpenStudy (experimentx):

first what kind of equation is that?? and how do you solve it??

OpenStudy (anonymous):

This is (I think) a first order seperable differential equation. You solve by having: \[f(x)dx = f(y)dy \] or

OpenStudy (anonymous):

Because you have to integrate both sides.

OpenStudy (anonymous):

SO you have to attach the y terms with dy, x terms with dx.

OpenStudy (experimentx):

it's called linear first order differential equation!!

OpenStudy (apoorvk):

Umm... you don't need that really. If a differential equation is of the form: \[\frac{dy}{dx} + Py = Q\] where both 'P' and 'Q' are functions of 'x', then, \[y.e^{\int P.dx}=\int Q.e^{\int P.dx}.dx ~+~k\] this is called a linear differential equation.

OpenStudy (apoorvk):

\[Here,~e^{\int P.dx}~is~known~as~the~INTEGRATING~FACTOR.\]

OpenStudy (anonymous):

I think I am going to read further about that. Right now, I would like you to help me with a question that I have already read about: \[\int\limits_{\pi/2}^{\pi} 34\csc(x)dx\] Is this convergent or divergent?

OpenStudy (anonymous):

I have got up to this part: \[\lim_{a \rightarrow \pi-} \ln|\csc(a) - \cot(a)|\] I then end up evaluating it and I get 0 for the argument in the ln.

OpenStudy (anonymous):

\[\int\limits\limits_{\pi/2}^{\pi} 34\csc(x)dx = \lim_{a \rightarrow \pi-} \int\limits\limits_{\pi/2}^{a} 34\csc(x)dx\]

OpenStudy (anonymous):

I know the integral of \[\csc(x)\]

OpenStudy (anonymous):

\[\int\limits_{}^{}\csc(x) dx = \ln|\csc(x) - \cot(x)| + C\]

OpenStudy (anonymous):

I think I can easily show that this integral is in fact convergent by the comparison theorem. However, I am having trouble showing this from the integral in question itself.

OpenStudy (anonymous):

@shivam_bhalla @amistre64 @satellite73 @FoolForMath @Zarkon @ParthKohli

OpenStudy (anonymous):

@TuringTest @Mertsj @Ishaan94

OpenStudy (experimentx):

try asking wolf

OpenStudy (anonymous):

No thanks. I need to show a fully justified solution. I KNOW that it converges, but I do not KNOW how to show it because I keep getting 0 in the argument of the natural log

OpenStudy (anonymous):

@Ishaan94 we are discussing not the differential equation, but the convergence question later on

OpenStudy (anonymous):

Oh Lool I was typing the solution to differential equation.

OpenStudy (anonymous):

@Mertsj

OpenStudy (anonymous):

@AccessDenied @FoolForMath @apoorvk @eliassaab

OpenStudy (experimentx):

\[ \int \tan x dx = - \ln |\cos x|\] Integrating Factor is \[ e^{-\ln|\cos x|} = \sec x\] your solution is \[ y = (\sec x) ^{- 1} \int \sec x 2 \cos^2x dx + C(\sec x)^{-1}\] \[ y = 2\cos x \int \cos x dx + C\cos x\] \[ y = 2\cos x (-\sin x) + C\cos x\] You know how to solve from here right??

OpenStudy (anonymous):

Sorry I was talking about the convergence question.

OpenStudy (anonymous):

Maybe I should post a new question for that. Hold on

OpenStudy (experimentx):

Oh ... LOL i guess you should ... i got confused